OpenStudy (anonymous):

find the x intercepts of a parabola with vertex (1,-108) and y intercept (0,-105)

OpenStudy (cwrw238):

long time since i done these...

OpenStudy (cwrw238):

general form is y = ax^2 + bx + c when x = 0 , y = -105 so c = -105

OpenStudy (cwrw238):

x coordinate of the vertex = -b / 2a = 1 the y coordinate of the vertex = -108 so -108 = a*1^2 + b - 105

OpenStudy (cwrw238):

so b = -2a and -108 = a + b - 105 a + b = -3 substitute b = -2a in the above: a - 2a = -3 -a = -3 a = 3 now finding b: b = -3 -3 = -6 so equation is 3x^2 - 6x - 105 = 0 solve to find the x intercepts: (3x - 21)(x + 5 ) = 0 x = 7 , -5 x intercepts are 7 and -5